Serial numbers for a product are to be made using 4 letters followed by 4numbersif the letters are to be taking from the first 6 letters of the alphabet with no repeat as and the numbers are taken from the the digit 0 through 9 with no repeats how many serial numbers can be generated
Answer by ikleyn(52781)

.
There are  6*5*4*3 = 360  4-letter words made of 6 letters without repetition,  and


There are  10*9*8*7 = 5040 4-digit numbers made of 10 digits from 0 to 9 without repetition.


In all, there are 650*5040 = 1814400 possible serial numbers.